OpenStudy (anonymous):

Please show step to step solution. Factorize (2014)^2 - (2013)^2

OpenStudy (loser66):

a^2 -b^2 = (a+b)(a-b) apply to your expression.

OpenStudy (anonymous):

Thank you for your help....

OpenStudy (loser66):

that means it doesn't help, right? you want the answer only, right? hehehe... then I am useless.

OpenStudy (anonymous):

No no, it really helps....

OpenStudy (anonymous):

I'm actually trying it myself.

OpenStudy (loser66):

:)

OpenStudy (anonymous):

Is the answer 4027, this is my work: (2014 + 2013) x (2014-2013) = 4027 x 1= 4027.